OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below. Show your work. x2 – 36 = 0 and x2 = 8x – 12 Part 2: Describe what the solution(s) represent to the graph of each. Part 3: How are the graphs alike? How are they different?

jhonyy9 (jhonyy9):

x2 -36=0 x2 = 36 x_1,_2= +/- sqrt36 = +/- 6 x2 -8x +12=0 (x-6)(x-2)=0 x-6=0 x=6 and x-2=0 x=2